# Plasticity and the corresponding mechanisms of nanopowdered Mg during shock consolidation and spallation

### 01 Mar 2024 v1

@@ I Introduction
Ultrafine grained (UFG) solids with grain size () of 1 m, and nanocrystalline (NC) materials with 100 nm (Chen et al., 2003; Hirth and Lothe, 1992), represent two distinct material types, owing to their unique physical properties, i.e., higher strength/hardness and toughness (Wang et al., 2002, 2006; Mo and Szlufarska, 2007), excellent diffusivity (Baró et al., 2001; Kim et al., 2000; Kolobov et al., 2001), and enhanced thermal expansion coefficients (Zhang et al., 1998). Such excellent properties and potentials prompt us to fabricate bulk UFG/NC materials, and two primary methods are developed, i.e., the top–down and bottom–up approaches (Ahn et al., 2015). For top–down approaches, the UFG/NC structures are produced via deforming grain refinement technique, i.e., the equal-channel angular pressing (ECAP) (Zhao et al., 2016; Langdon, 2007; Quang et al., 2007) and high-pressure rolling (HPT) (Horita and Langdon, 2005; Edalati and Horita, 2016; Azzeddine et al., 2022). For bottom–up approaches, UFG/NC materials are manufactured via powder metallurgy, such as inert gas condensation (Gleiter, 1989; Zheng and Branicio, 2020; Patelli et al., 2022), spark plasma sintering (Hu et al., 2020; Olevsky et al., 2007; Cavaliere, 2019), and high pressure cold sintering (Guo et al., 2019).
@@
Shock consolidation (Schwarz et al., 1984; Zhang et al., 2008; Gourdin, 1986), is a unique bottom–up method, by compacting nanopowders (NP) to strong bulk NC materials using shock waves from an explosive or high-speed collision (Jin et al., 2004; Chen et al., 2004; Zhang et al., 2022). When shock waves pass through, it contributes to an ultra-high pressure in the powders. Then the particle undergoes an apparent particle deformation and densification, due to a collapse of powder agglomerates (Zhou and Chen, 2016). Compared to conventional methods, shock consolidation is one of the most efficient methods to manufacture bulk materials with improved properties (Meyers, 1994; Thadhani, 1993; Viswanathan et al., 2006). For instance, the structure of the starting powders retains and no grain growth occurs during shock consolidation (Meyers et al., 1999), due to extraordinarily fast process within a microsecond (s) time scale. Consequently, lots of bulk metallic/alloy nanostructured materials, such as aluminum (Nieh et al., 1996; Venz et al., 2003; Brochu et al., 2007; Fredenburg et al., 2010), copper (Huang et al., 2012), iron (Dai et al., 2008), silver (Zhang et al., 2008), tungsten–copper (Zhou and Chen, 2016; Wang et al., 2010), high-entropy alloys (Yim et al., 2017), and aluminum ceramic (Bukaemskii and Fedorova, 2008), are fabricated using shock consolidations. However, the details of consolidation process, are rarely unveiled until now, which determine the structures and properties of final bulk nanostructured materials (Feng et al., 2020). In facts, capturing the microstructural features in dynamics and the corresponding mechanisms, is a key to optimize and control the dynamic powder compaction process.
@@
Recently, both experiment and simulation efforts are devoted to understanding the mechanisms of shock consolidation. Experimentally, Matsumoto (Matsumoto and Kondo, 1989) found that molten jet, dynamic friction of particles, and the plastic deformation around a void, facilitate the compaction, using an analysis of scanning electronic microscope (SEM). It is validated by the following experiments on fabrication of fine-grained W-Cu composites (Zhou and Chen, 2016). In Kondo’s experiment (Kondo and Sawai, 1990), it presented that local heating, induced by plasticity, accelerates the consolidation. Gao (Gao et al., 1991) proposed a melting-welding consolidation mechanism based on the experiment for metal powder with a random arrangement. However, the technique difficulties for current set-ups, in probing the structural deformation on the atomic scale and ultrashort time scale (Liu et al., 2019b), limit the knowledge on the dynamic process during shock impact in depth. Recent developed femtosecond x-ray diffraction (XRD) using x-ray free-electron laser (XFEL) (Wenk et al., 2013), provides opportunities for determining the irreversible transient processes of plasticity deformation. However, we are not aware of such experiments in shock consolidation, although they have potentials for real-time, , and high-resolution temporal and spatial probes (Chen et al., 2019; Chai et al., 2020; Turneaure et al., 2020).
@@
Theoretically, the continuum models are implemented, to describe the compaction of powders in dynamics. They can help in both understanding the experiments and accessing the regimes that inaccessible in current experiments (Rodriguez-Nieva et al., 2014). In Boltachev’s model (Boltachev et al., 2009) for granular medium, the evolution of boundaries between powder and the container, are taken into account in compaction model to describe the the mechanical response quantitatively. Ahn . (Ahn et al., 2015) then developed a Cu-compacted model by incorporating a dislocation model. They also found a strong dependence of dislocation for the propagation of shock wave and the void-collapse induced plastic flow. These models are appropriated to describ the deformation and its underlying mechanisms at mesoscale, but the validity of models in microscopic is not clear. For understanding plasticity and defect procedures of nanopowders during consolidation, at the microscopic level, MD simulations are employed. Huang . (Huang et al., 2012) studied shock consolidation of Cu nanopowders and the dynamic tensile response in their simulations. It is found that dislocation, lattice rotation, shearing, and friction contribute to the void collapse. Mayer et al. (Mayer et al., 2020) reported a strong relation between dislocation slips and compaction in aluminum nanopowder, and proposed a mechanical model for describing metal nanopowder compaction. Feng . (Feng et al., 2020) investigated the shock compaction of tungsten nanoparticles, showing a liquid-diffusion/solid-pressure welding mechanism during consolidation.
@@
In this work, we systematically investigate the shock response and its microstructure deformation of magnesium (Mg) nanopowders during shock consolidation, by conducting nonequilibrium molecular dynamics (NEMD) simulations (Wang et al., 2019) and XRD analysis. Here, Mg has been chosen mainly because of its wide applications in aerospace, automotive, medical equipment, national defense science and industry (Chen et al., 2015; Lentz et al., 2016; Liu et al., 2019a; Xu et al., 2018). This paper is organized as follows: a description of the details of MD simulations and related analysis methodology is presented in Sec.II, the results of simulations and discussion in Sec. III, followed by summary and conclusions in Sec. IV.
@@ II Methodology
For our NEMD simulations, the open source code Large scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) (Plimpton, 1995), and modified embedded atom method (MEAM) potential of Mg (Kim et al., 2009), are utilized. This potential can accurately describe the plasticity of Mg (Yi et al., 2016; Tang et al., 2023), verified by density function theory (Ghazisaeidi et al., 2014; Wang and Liu, 2022) and experimental results (Sun et al., 2018). This potential also presents a reasonable accuracy for shock simulations. To check the performance of such MEAM potential during shock loading, we carry out simulations under different shock strengths in single- and nano-crystal Mg. For describing the effect of crystallographic orientation in single crystal, we choose the -axis, parallel to the [110], [100], or [0001] directions as the impact direction. The dimensions of the single-crystal Mg are about 145 13 13 nm ( 1.2 million atoms), and the impact velocities range from 1.0 to 4.0 km s. Based on the – relations of the plastic shocks (Fig. 1(a)), all simulation results, including single-crystals (shocked along [110], [100], and [0001], respectively), nanocrystals, and nanopowders, are in agreement with the experimental results (Baytos, 1980). The functions of pressure vs normalized specific volume (/) are shown in Fig. 1(b). Despite the scatter in the experimental data, our simulation results are consistent with the experiments for the compression (1-/), ranging from 6 to 44. It presents a reasonable accuracy of this potential for shock simulations from the results above, although its accuracy for deformation at high pressure ( 4.0 km s) remains to be established.
@@
We construct an idealized columnar nanopowdered configuration, containing 18 grains of identical shape and diameter ( 18 nm) with four different crystallographic orientations (A–D, Fig. 2), in the three-dimensional (3D) periodic cell. The columnar axis is along [100] crystallographic direction (the -axis), and the thickness along this axis is about 30 nm. The grains are oriented relative to the -axis ([110] in grain A), by an angle of = 0, 30, 60, and 90 for grains A–D, respectively. The dimensions of the configurations are about 156 30 30 nm, and containing approximately 5.6 million atoms. The configurations are first relaxed at 0 K with the conjugate gradient method, and then thermalized at ambient conditions with a constant-pressure-temperature ensemble and 3D periodic boundary conditions, prior to shock loading. Shock simulations are then performed with the microcanonical ensemble. Periodic boundary conditions are applied along the - and -axes, while a free boundary is applied along the -axis. The timestep for integration of the equation of motion is 1 fs. The small dense region on the left is set as the piston (Holian and Lomdahl, 1998) in our shock simulations. The interactions between the piston and the reset of the atoms in the configuration are described with the same interatomic potential, while the atoms in the piston do not participate in molecular dynamics. The atomic piston delivers the shock with the piston velocity of , ranging from 0.5 to 3.0 km s, starting at time = 0 from = 0 along the -axis towards the free surface. At a certain time (), the piston is stopped, creating a release fan propagating towards the free surface. The release fan then tend to interact with each other, due to shock reflection at the free surface, and gives rise to release and subsequent tension within the target interior (Luo et al., 2009).
@@
The one-dimensional (1D) and two-dimensional (2D) binning analyses (Huang et al., 2012), resolving spatially physical properties such as stress (, = , , and ), are performed to describe the response during shock loading. The bin width is 0.5 nm. The center-of-mass velocity of a bin is removed when calculating within each bin: = -(/), where is the atomic mass and is the atomic volume averaged over the bin. Then the pressure and shear stress can be calculated (Wang et al., 2019). To characterize the structural deformation, the common neighbor analysis (CNA) (Tsuzuki et al., 2007; Ackland and Jones, 2006) and slip vector analysis (Wang et al., 2014), are implemented. To better visualize the plastic deformation in dynamics, the orientation mapping (OM) analysis, following the standard electron backscatter diffraction (EBSD) analysis, is implemented as a complement to atomic-level characterization. The simulated the X-ray diffraction (XRD) patterns (Wang et al., 2015; Bi et al., 2023) are also calculated, to revel the the microstructure features and deformation in real time, by performing a parallel code of GAPD (E et al., 2018). Such XRD analysis is widely applied to quantitatively determine the microstructure characteristics (Chen et al., 2019; Zhang et al., 2021; Mo et al., 2022; Zhang et al., 2023), by using free electron lasers and synchrotron radiation in experimients.
@@ III.1.1 Shock consolidation: “heterogeneous” plasticity and “homogeneous” disordering
Upon impact, the shock wave along the -direction prompts an apparent consolidation in NP-Mg, contributed by both plastic deformation and pore collapse. Two different structural deformation modes are included during shock compression: i) the stacking faults (SF) and deformation twinning at lower impact velocities ( 1.0 km s; Figs. 3(a) and (b)), conducing to the grain rotation and subgrains formation, and ii) the disordering at higher impact velocities ( 2.0 km s; Figs. 3(c) and (d)). Such disordering consist of the amorphization (or solid disordering) at = 2.0 km s, and the solid-liquid transition (or shock-induced melting) at = 3.0 km s. The calculated diffusion coefficients () are 10–10 m s for the former, similar to the solid crystal ( 10 m s); whereas 10 m s for the latter, approaching the liquid state ( 10 m s) (He et al., 2014).
@@ III.1.2 Deformation twinning
Deformation twinning, prompting the grain rotation partially, i.e., (parent) (variant), is a primary deformation mode in NP-Mg during shock compression as 1.0 km s (Fig. 3(a) and (b)). Two extension twinning modes, i.e., the (T) (Zhou et al., 2019; Flanagan et al., 2020), and twins (T) (Wang et al., 2020; Zhang et al., 2020), and a compression twinning mode, i.e., the (T) (Zhou et al., 2019) twin, as the secondary twin following the primary twin (T), are activated in NP-Mg, which also present apparent crystallographic orientation dependence.
A. twinning. When 1.0 km s, the twinning is the primary deformation mode in A- and B-type grains, as shown in Fig. 3(a) and (b). It contributes to a grain rotation with a rotation angle, , (the misorientation angle between the parent and variant) of 33.2 around axis, and induces a transition (Fig. 6(a) and (b)). It is manifested via analyzing the diffraction pattern (Fig. 6(c)), with the development of high temporal and spatial resolution x-ray probes. As twinning proceeds, diffraction spots tend to be diffused and separated in diffraction pattern (in view of the plane). Two characteristics of twinning are presented: (i) the rotation angle between and pattern is about 32.7, consistent with the one determined in real space. The rotation angle is well agreement with the experiments results ( 31–35) for twins (Zhou et al., 2019; Tang et al., 2023; Christian and Mahajan, 1995); (ii) the diffraction spots from the reciprocal space, i.e., 111- and 11-, are coincident with each other, via analyzing the features in patterns, implying the parent () and twin variant () have the common symmetric plane . For twin, its geometric relation can be presented in Fig. 6(d), i.e., the twinning plane (111) (dark blue) and the corresponding slip direction (26) (red arrow).
What is the mechanism governing such deformation twinning? The fact is that such twinning is prompted by coupling the activation of slip systems (Fig. 6(e)) and atomic shuffles (Ishii et al., 2016) by atoms moving along the [100] or [100] directions in the planes (Fig. 6(f)), based on the analysis of slip vectors. The atomic slips, achieved via the consecutive emission of [26]/6 partials on neighboring (111) planes, facilitate the rotations from the parent hcp () to the variant () unit cell by 33 (Fig. 6(g)), via shifting the stacking sequence from ABAB to ACAC; the subsequent lattice reorientation via atomic shuffles in variant regions on (0001) planes (Fig. 6(h)), contribute to the mirroring structure about the (111) plane, mediating a shift from the incoherent prismatic boundary (IPB) to the coherent twin boundary (CTB: (111) plane).
@@
What is the underlying mechanism of such double-twin, and what is the relation between the primary and secondary twinning? Essentially, the coupling between atomic shuffles and slips conduces to the double-twinning, similar to (111) deformation twinning, based on the slip vector analysis (Fig. 7(e) and (f)). Considering the initial configuration as the reference, similar to the postmortem analysis, it presents that the primary is driven by slips and atomic shuffles, along directions, while is prompted via the slips ( system). When the intermediate configuration (only variant is activated within grains) is considered as the reference, is governed by atomic shuffles and slips, similar to that for (insets, Fig. 7(e) and (f)). It presents the different mechanisms for secondary twinning, i.e., , based on the analysis above. Such difference can be attributed to the different analysis modes. For the former, the conclusions are obtained via directly comparing the initial and final configuration, while the dynamic process, i.e., the precursor, which accelerates the activation of via reducing its energy barrier (Peng et al., 2015), is ignored. Then we duduce the procedures of double twinning based on the discussion above (Fig. 7(g)): i) prompted by atomic shuffles and slips, it first facilitates the emission and growth of , along with the primary twinning; and ii) inside variant, another twinning is activated, which conduces to the rearrangement of atomic configurations, i.e., , and the formation of twin boundary (TB) between and . With the increase of secondary variant, i.e., , towards the parent, it then triggers a transition of TB from the to between the parent () and variants ( for the former, and for the latter). Simultaneously, the misorientation angle between the parent () and variant changes from 33 to 61, owing to the lattice reorientation within the variant.
@@
In previous studies (Wang et al., 2019), deformation twinning is considered as an intermediate state, which contributes to the lattice reorientation, as the final state, during shock impact. However, twinning with misorientation angle of 86.3, can be frequently observed in HCP metals, via performing the postmorten microstructural analysis using SEM and TEM (Bhattacharyya et al., 2009; Zhang et al., 2012; Yu et al., 2014; Zhou et al., 2019). It implies that the twinning is a stable rather than transient phase during shock-induced plasticity. Furthermore, the discussion above also presents a sequence in shock-induced deformation in HCP-Mg, i.e., the preferential lattice reorientation and the following twinning. Consequently, it is then deduced that deformation twinning is contributed by coupling the transient lattice reorientation and slips. Upon impact, atomic shuffles, by atoms moving along directions in the planes, based on the analysis of slip vectors (Fig. 8(e) and (f)), first prompt the apparent rotations of the parent by 90, around direction (Fig. 8(g)); and the subsequent emission of (102)[101] slips facilitates the lattice reorientation to deformation twinning, with a misorientation angle of 86.3 (Fig. 8(g)).

### 13 Mar 2024 v2

@@ I Introduction
Ultrafine grained (UFG) solids with grain size () of 1 m, and nanocrystalline (NC) materials with 100 nm Chen et al. (2003); Hirth and Lothe (1992), represent two distinct material types, owing to their unique physical properties, i.e., higher strength/hardness and toughness Wang et al. (2002, 2006); Mo and Szlufarska (2007), excellent diffusivity Baró et al. (2001); Kim et al. (2000); Kolobov et al. (2001), and enhanced thermal expansion coefficients Zhang et al. (1998). Such excellent properties and potentials prompt us to fabricate bulk UFG/NC materials, and two primary methods are developed, i.e., the top–down and bottom–up approaches Ahn et al. (2015). For top–down approaches, the UFG/NC structures are produced via deforming grain refinement technique, i.e., the equal-channel angular pressing (ECAP) Zhao et al. (2016); Langdon (2007); Quang et al. (2007) and high-pressure rolling (HPT) Horita and Langdon (2005); Edalati and Horita (2016); Azzeddine et al. (2022). For bottom–up approaches, UFG/NC materials are manufactured via powder metallurgy, such as inert gas condensation Gleiter (1989); Zheng and Branicio (2020); Patelli et al. (2022), spark plasma sintering Hu et al. (2020); Olevsky et al. (2007); Cavaliere (2019), and high pressure cold sintering Guo et al. (2019).
@@
Shock consolidation Schwarz et al. (1984); Zhang et al. (2008); Gourdin (1986), is a unique bottom–up method, by compacting nanopowders (NP) to strong bulk NC materials using shock waves from an explosive or high-speed collision Jin et al. (2004); Chen et al. (2004); Zhang et al. (2022). When shock waves pass through, it contributes to an ultra-high pressure in the powders. Then the particle undergoes an apparent particle deformation and densification, due to a collapse of powder agglomerates Zhou and Chen (2016). Compared to conventional methods, shock consolidation is one of the most efficient methods to manufacture bulk materials with improved properties Meyers (1994); Thadhani (1993); Viswanathan et al. (2006). For instance, the structure of the starting powders retains and no grain growth occurs during shock consolidation citemeyers99am, due to extraordinarily fast process within a microsecond (s) time scale. Consequently, lots of bulk metallic/alloy nanostructured materials, such as aluminum Nieh et al. (1996); Venz et al. (2003); Brochu et al. (2007); Fredenburg et al. (2010), copper Huang et al. (2012), iron Dai et al. (2008), silver Zhang et al. (2008), tungsten–copper Zhou and Chen (2016); Wang et al. (2010), high-entropy alloys Yim et al. (2017), and aluminum ceramic Bukaemskii and Fedorova (2008), are fabricated using shock consolidations. However, the details of consolidation process, are rarely unveiled until now, which determine the structures and properties of final bulk nanostructured materials Feng et al. (2020). In facts, capturing the microstructural features in dynamics and the corresponding mechanisms, is a key to optimize and control the dynamic powder compaction process.
@@
Recently, both experiment and simulation efforts are devoted to understanding the mechanisms of shock consolidation. Experimentally, Matsumoto Matsumoto and Kondo (1989) found that molten jet, dynamic friction of particles, and the plastic deformation around a void, facilitate the compaction, using an analysis of scanning electronic microscope (SEM). It is validated by the following experiments on fabrication of fine-grained W-Cu composites Zhou and Chen (2016). In Kondo’s experiment Kondo and Sawai (1990), it presented that local heating, induced by plasticity, accelerates the consolidation. Gao Gao et al. (1991) proposed a melting-welding consolidation mechanism based on the experiment for metal powder with a random arrangement. However, the technique difficulties for current set-ups, in probing the structural deformation on the atomic scale and ultrashort time scale Liu et al. (2019a), limit the knowledge on the dynamic process during shock impact in depth. Recent developed femtosecond x-ray diffraction (XRD) using x-ray free-electron laser (XFEL) Wenk et al. (2013), provides opportunities for determining the irreversible transient processes of plasticity deformation. However, we are not aware of such experiments in shock consolidation, although they have potentials for real-time, , and high-resolution temporal and spatial probes Chen et al. (2019); Chai et al. (2020); Turneaure et al. (2020).
@@
Theoretically, the continuum models are implemented, to describe the compaction of powders in dynamics. They can help in both understanding the experiments and accessing the regimes that inaccessible in current experiments Rodriguez-Nieva et al. (2014). In Boltachev’s model Boltachev et al. (2009) for granular medium, the evolution of boundaries between powder and the container, are taken into account in compaction model to describe the the mechanical response quantitatively. Ahn . Ahn et al. (2015) then developed a Cu-compacted model by incorporating a dislocation model. They also found a strong dependence of dislocation for the propagation of shock wave and the void-collapse induced plastic flow. These models are appropriated to describ the deformation and its underlying mechanisms at mesoscale, but the validity of models in microscopic is not clear. For understanding plasticity and defect procedures of nanopowders during consolidation, at the microscopic level, MD simulations are employed. Huang . Huang et al. (2012) studied shock consolidation of Cu nanopowders and the dynamic tensile response in their simulations. It is found that dislocation, lattice rotation, shearing, and friction contribute to the void collapse. Mayer et al. Mayer et al. (2020) reported a strong relation between dislocation slips and compaction in aluminum nanopowder, and proposed a mechanical model for describing metal nanopowder compaction. Feng . Feng et al. (2020) investigated the shock compaction of tungsten nanoparticles, showing a liquid-diffusion/solid-pressure welding mechanism during consolidation.
@@
In this work, we systematically investigate the shock response and its microstructure deformation of magnesium (Mg) nanopowders during shock consolidation, by conducting nonequilibrium molecular dynamics (NEMD) simulations Wang et al. (2019) and XRD analysis. Here, Mg has been chosen mainly because of its wide applications in aerospace, automotive, medical equipment, national defense science and industry Chen et al. (2015); Lentz et al. (2016); Liu et al. (2019b); Xu et al. (2018). This paper is organized as follows: a description of the details of MD simulations and related analysis methodology is presented in Sec.II, the results of simulations and discussion in Sec. III, followed by summary and conclusions in Sec. IV.
@@ II Methodology
For our NEMD simulations, the open source code Large scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) Plimpton (1995), and modified embedded atom method (MEAM) potential of Mg Kim et al. (2009), are utilized. This potential can accurately describe the plasticity of Mg Yi et al. (2016); Tang et al. (2023), verified by density function theory Ghazisaeidi et al. (2014); Wang and Liu (2022) and experimental results Sun et al. (2018). This potential also presents a reasonable accuracy for shock simulations. To check the performance of such MEAM potential during shock loading, we carry out simulations under different shock strengths in single- and nano-crystal Mg. For describing the effect of crystallographic orientation in single crystal, we choose the -axis, parallel to the [110], [100], or [0001] directions as the impact direction. The dimensions of the single-crystal Mg are about 145 13 13 nm ( 1.2 million atoms), and the impact velocities range from 1.0 to 4.0 km s. Based on the – relations of the plastic shocks (Fig. 1(a)), all simulation results, including single-crystals (shocked along [110], [100], and [0001], respectively), nanocrystals, and nanopowders, are in agreement with the experimental results Baytos (1980). The functions of pressure vs normalized specific volume (/) are shown in Fig. 1(b). Despite the scatter in the experimental data, our simulation results are consistent with the experiments for the compression (1-/), ranging from 6 to 44. It presents a reasonable accuracy of this potential for shock simulations from the results above, although its accuracy for deformation at high pressure ( 4.0 km s) remains to be established.
@@
We construct an idealized columnar nanopowdered configuration, containing 18 grains of identical shape and diameter ( 18 nm) with four different crystallographic orientations (A–D, Fig. 2), in the three-dimensional (3D) periodic cell. The columnar axis is along [100] crystallographic direction (the -axis), and the thickness along this axis is about 30 nm. The grains are oriented relative to the -axis ([110] in grain A), by an angle of = 0, 30, 60, and 90 for grains A–D, respectively. The dimensions of the configurations are about 156 30 30 nm, and containing approximately 5.6 million atoms. The configurations are first relaxed at 0 K with the conjugate gradient method, and then thermalized at ambient conditions with a constant-pressure-temperature ensemble and 3D periodic boundary conditions, prior to shock loading. Shock simulations are then performed with the microcanonical ensemble. Periodic boundary conditions are applied along the - and -axes, while a free boundary is applied along the -axis. The timestep for integration of the equation of motion is 1 fs. The small dense region on the left is set as the piston Holian and Lomdahl (1998) in our shock simulations. The interactions between the piston and the reset of the atoms in the configuration are described with the same interatomic potential, while the atoms in the piston do not participate in molecular dynamics. The atomic piston delivers the shock with the piston velocity of , ranging from 0.5 to 3.0 km s, starting at time = 0 from = 0 along the -axis towards the free surface. At a certain time (), the piston is stopped, creating a release fan propagating towards the free surface. The release fan then tend to interact with each other, due to shock reflection at the free surface, and gives rise to release and subsequent tension within the target interior Luo et al. (2009).
@@
The one-dimensional (1D) and two-dimensional (2D) binning analyses Huang et al. (2012), resolving spatially physical properties such as stress (, = , , and ), are performed to describe the response during shock loading. The bin width is 0.5 nm. The center-of-mass velocity of a bin is removed when calculating within each bin: = -(/), where is the atomic mass and is the atomic volume averaged over the bin. Then the pressure and shear stress can be calculated Wang et al. (2019). To characterize the structural deformation, the common neighbor analysis (CNA) Tsuzuki et al. (2007); Ackland and Jones (2006) and slip vector analysis Wang et al. (2014), are implemented. To better visualize the plastic deformation in dynamics, the orientation mapping (OM) analysis, following the standard electron backscatter diffraction (EBSD) analysis, is implemented as a complement to atomic-level characterization. The simulated the X-ray diffraction (XRD) patterns Wang et al. (2015); Bi et al. (2023) are also calculated, to revel the the microstructure features and deformation in real time, by performing a parallel code of GAPD E et al. (2018). Such XRD analysis is widely applied to quantitatively determine the microstructure characteristics Chen et al. (2019); Zhang et al. (2021); Mo et al. (2022); Zhang et al. (2023), by using free electron lasers and synchrotron radiation in experimients.
@@ III.1.1 Shock consolidation: “heterogeneous” plasticity and “homogeneous” disordering
Upon impact, the shock wave along the -direction prompts an apparent consolidation in NP-Mg, contributed by both plastic deformation and pore collapse. Two different structural deformation modes are included during shock compression: i) the stacking faults (SF) and deformation twinning at lower impact velocities ( 1.0 km s; Figs. 3(a) and (b)), conducing to the grain rotation and subgrains formation, and ii) the disordering at higher impact velocities ( 2.0 km s; Figs. 3(c) and (d)). Such disordering consist of the amorphization (or solid disordering) at = 2.0 km s, and the solid-liquid transition (or shock-induced melting) at = 3.0 km s. The calculated diffusion coefficients () are 10–10 m s for the former, similar to the solid crystal ( 10 m s); whereas 10 m s for the latter, approaching the liquid state ( 10 m s) He et al. (2014).
@@ III.1.2 Deformation twinning
Deformation twinning, prompting the grain rotation partially, i.e., (parent) (variant), is a primary deformation mode in NP-Mg during shock compression as 1.0 km s (Fig. 3(a) and (b)). Two extension twinning modes, i.e., the (T) Zhou et al. (2019); Flanagan et al. (2020), and twins (T) Wang et al. (2020); Zhang et al. (2020), and a compression twinning mode, i.e., the (T) Zhou et al. (2019) twin, as the secondary twin following the primary twin (T), are activated in NP-Mg, which also present apparent crystallographic orientation dependence.
A. twinning. When 1.0 km s, the twinning is the primary deformation mode in A- and B-type grains, as shown in Fig. 3(a) and (b). It contributes to a grain rotation with a rotation angle, , (the misorientation angle between the parent and variant) of 33.2 around axis, and induces a transition (Fig. 6(a) and (b)). It is manifested via analyzing the diffraction pattern (Fig. 6(c)), with the development of high temporal and spatial resolution x-ray probes. As twinning proceeds, diffraction spots tend to be diffused and separated in diffraction pattern (in view of the plane). Two characteristics of twinning are presented: (i) the rotation angle between and pattern is about 32.7, consistent with the one determined in real space. The rotation angle is well agreement with the experiments results ( 31–35) for twins Zhou et al. (2019); Tang et al. (2023); Christian and Mahajan (1995); (ii) the diffraction spots from the reciprocal space, i.e., 111- and 11-, are coincident with each other, via analyzing the features in patterns, implying the parent () and twin variant () have the common symmetric plane . For twin, its geometric relation can be presented in Fig. 6(d), i.e., the twinning plane (111) (dark blue) and the corresponding slip direction (26) (red arrow).
What is the mechanism governing such deformation twinning? The fact is that such twinning is prompted by coupling the activation of slip systems (Fig. 6(e)) and atomic shuffles Ishii et al. (2016) by atoms moving along the [100] or [100] directions in the planes (Fig. 6(f)), based on the analysis of slip vectors. The atomic slips, achieved via the consecutive emission of [26]/6 partials on neighboring (111) planes, facilitate the rotations from the parent hcp () to the variant () unit cell by 33 (Fig. 6(g)), via shifting the stacking sequence from ABAB to ACAC; the subsequent lattice reorientation via atomic shuffles in variant regions on (0001) planes (Fig. 6(h)), contribute to the mirroring structure about the (111) plane, mediating a shift from the incoherent prismatic boundary (IPB) to the coherent twin boundary (CTB: (111) plane).
@@
What is the underlying mechanism of such double-twin, and what is the relation between the primary and secondary twinning? Essentially, the coupling between atomic shuffles and slips conduces to the double-twinning, similar to (111) deformation twinning, based on the slip vector analysis (Fig. 7(e) and (f)). Considering the initial configuration as the reference, similar to the postmortem analysis, it presents that the primary is driven by slips and atomic shuffles, along directions, while is prompted via the slips ( system). When the intermediate configuration (only variant is activated within grains) is considered as the reference, is governed by atomic shuffles and slips, similar to that for (insets, Fig. 7(e) and (f)). It presents the different mechanisms for secondary twinning, i.e., , based on the analysis above. Such difference can be attributed to the different analysis modes. For the former, the conclusions are obtained via directly comparing the initial and final configuration, while the dynamic process, i.e., the precursor, which accelerates the activation of via reducing its energy barrier Peng et al. (2015), is ignored. Then we duduce the procedures of double twinning based on the discussion above (Fig. 7(g)): i) prompted by atomic shuffles and slips, it first facilitates the emission and growth of , along with the primary twinning; and ii) inside variant, another twinning is activated, which conduces to the rearrangement of atomic configurations, i.e., , and the formation of twin boundary (TB) between and . With the increase of secondary variant, i.e., , towards the parent, it then triggers a transition of TB from the to between the parent () and variants ( for the former, and for the latter). Simultaneously, the misorientation angle between the parent () and variant changes from 33 to 61, owing to the lattice reorientation within the variant.
@@
In previous studies Wang et al. (2019), deformation twinning is considered as an intermediate state, which contributes to the lattice reorientation, as the final state, during shock impact. However, twinning with misorientation angle of 86.3, can be frequently observed in HCP metals, via performing the postmorten microstructural analysis using SEM and TEM Bhattacharyya et al. (2009); Zhang et al. (2012); Yu et al. (2014); Zhou et al. (2019). It implies that the twinning is a stable rather than transient phase during shock-induced plasticity. Furthermore, the discussion above also presents a sequence in shock-induced deformation in HCP-Mg, i.e., the preferential lattice reorientation and the following twinning. Consequently, it is then deduced that deformation twinning is contributed by coupling the transient lattice reorientation and slips. Upon impact, atomic shuffles, by atoms moving along directions in the planes, based on the analysis of slip vectors (Fig. 8(e) and (f)), first prompt the apparent rotations of the parent by 90, around direction (Fig. 8(g)); and the subsequent emission of (102)[101] slips facilitates the lattice reorientation to deformation twinning, with a misorientation angle of 86.3 (Fig. 8(g)).